# Fast Hip Joint Moment Estimation with A General Moment Feature Generation Method

### v1

@@ II-C GMF Estimator
Existing recurrent neural network structures include RNN, LSTM, GRU [29]. Due to the challenges of RNN in handling long-term dependencies, vanishing gradients, and exploding gradients, they have been gradually replaced by LSTM and GRU. Furthermore, GRU has a more simplified structure and fewer parameters compared to LSTM, resulting in lower memory usage and faster execution. In many cases, there is no significant difference in prediction accuracy between the two models [30]. Therefore, the estimator in this study utilizes a neural network based on the GRU structure.
@@
In this study, only one GRU block is used as an implementation case, with 16 hidden neurons, and a fully connected network with a size of 16*1 is used to output the predicted target (Fig. 2). Specifically, at each time step, the GRU receives two different modalities: the hidden information from the previous time step’s GRU hidden layer and the inputs $\boldsymbol{X}$. It adjusts the information flow through update gates and reset gates, and finally obtains the output based on these two gates. Thus, given the hip joint sagittal plane angle signal at time t as $\boldsymbol{X}_{t}$, the hidden inputs from the previous time step as $\boldsymbol{H}_{t-1}$, the estimated GMF at the current time step $\widehat{GMF}$ can be obtained as follows:
@@
$\boldsymbol{R}_{t}=\sigma(\boldsymbol{W}_{xr}\boldsymbol{X}_{t}+\boldsymbol{W}_{hr}\boldsymbol{H}_{t-1}+\boldsymbol{b}_{r})$ (8)
$\boldsymbol{Z}_{t}=\sigma(\boldsymbol{W}_{xz}\boldsymbol{X}_{t}+\boldsymbol{W}_{hz}\boldsymbol{H}_{t-1}+\boldsymbol{b}_{z})$ (9)
$\boldsymbol{C}_{t}=\text{tanh}(\boldsymbol{W}_{xc}\boldsymbol{X}_{t}+\boldsymbol{W}_{hc}(\boldsymbol{R}_{t}\odot\boldsymbol{H}_{t-1})+\boldsymbol{b}_{c})$ (10)
$\boldsymbol{H}_{t}=(1-\boldsymbol{Z}_{t})\odot\boldsymbol{H}_{t-1}+\boldsymbol{Z}_{t}\odot\boldsymbol{C}_{t}$ (11)
$\widehat{GMF}=\boldsymbol{W}_{FC}\boldsymbol{H}_{t}+\boldsymbol{b}_{FC},$ (12)
where $\boldsymbol{W}_{xr}$, $\boldsymbol{W}_{hr}$, $\boldsymbol{W}_{xz}$, $\boldsymbol{W}_{hz}$, $\boldsymbol{W}_{xc}$ and $\boldsymbol{W}_{hc}$ are the GRU layer weights, $\boldsymbol{b}_{r}$, $\boldsymbol{b}_{z}$ and $\boldsymbol{b}_{c}$ are biases, $\boldsymbol{R}_{t}$ and $\boldsymbol{Z}_{t}$ denote vectors for the activation values of the update gate and reset gate, and $\boldsymbol{W}_{fc}$ is the full connected layer weights and $\boldsymbol{b}_{fc}$ is the biases. Writing the network parameters ($\boldsymbol{W}$ and $\boldsymbol{b}$) as $\boldsymbol{\theta}_{\boldsymbol{E}}$, like (5), the optimal parameters $\boldsymbol{\theta}_{\boldsymbol{E}}^{*}$ are solved as:
$\boldsymbol{\theta}_{\boldsymbol{E}}^{*}=\mathop{\textbf{argmin}}\limits_{\boldsymbol{\theta}_{\boldsymbol{E}}}\left\|\boldsymbol{E}(\boldsymbol{X};\boldsymbol{\theta}_{\boldsymbol{E}})-\boldsymbol{G}(\boldsymbol{Q},M;\boldsymbol{\theta}_{\boldsymbol{G}})\right\|_{2}.$ (13)
II-D Model Training
The proposed method consists of two parts: GMF estimation and GMF decoding, which involves the training of three models: estimator, generator, and decoder. In the first part, the GMF generator is jointly trained with the estimator to obtain better individual-independent representation (GMF) and its prediction model (estimator). Meanwhile, in the second part, the GMF generator is trained together with the decoder to ensure the decodability of the generalized GMF and to obtain a model capable of decoding GMF into joint moment.
Based on (5), (7), and (13), two loss functions $L_{1}$, $L_{2}$ are constructed for the network. The backpropagation process is shown in Fig. 1 - Model training with solid lines. $L_{1}$ represents the loss between the results of the GMF estimator and the GMF generator, while $L_{2}$ represents the loss between the outputs of the GMF decoder and ground-truth values of joint moment.
$J_{mse}(y_{i},\hat{y_{i}})=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y_{i}})^{2}$ (14)
$L_{1}=J_{mse}(\boldsymbol{F}(\boldsymbol{X}),\boldsymbol{G}(\boldsymbol{Q},M))$ (15)
$L_{2}=J_{mse}(\boldsymbol{R}(\boldsymbol{G}(\boldsymbol{Q},M),\boldsymbol{Q}),M).$ (16)
It is worth noting that the optimization of the parameters of the GMF generator are included in $L_{2}$. By selecting the weighting coefficients $w_{1}$ and $w_{2}$ of the two loss functions, it can be ensured that the GMF, which is generated by the feature generator optimized by $L_{1}$, can be decoded into joint moment through the GMF decoder. Therefore, the condition in (5) is satisfied. The overall loss function of the joint moment estimation network is given by:
@@ II-D Model Training
II-E Model Prediction
The trained GMF estimator and decoder will be used in the prediction process for the final joint moment estimation. As shown in Fig. 1 - Prediction, using the joint kinematics as input, the GMF is obtained by the fast GMF estimator, which is then combined with user body parameters to be decoded into the predicted joint moment:
$\widehat{M}=\boldsymbol{R}(\boldsymbol{E}(\boldsymbol{X}),\boldsymbol{Q})$ (18)
III Experiment
III-A Dataset and Training-testing Data Partitioning
Figure 4: (a) Trend of accuracy of the joint moment prediction(dark blue) and accuracy of GMF decoding (light blue) of training data during the training process of the proposed estimation model based on GMF. Both values of them continuously improve and converge to the optimum over the 5000 epochs. (b) Trend of prediction accuracy on the validation set for our method (dark red) compared to two other contrastive methods (light red). The three methods approximate the best accuracy achieved under the experimental conditions set.
@@ III-A Dataset and Training-testing Data Partitioning
III-B Implementation Details
We conducted model training on a device with an Intel CPU (Core i7-12700F) and an Nvidia GPU (GeForce RTX 3060Ti). The Adam optimizer [32] was employed. In all experiments, we performed five repetitions with different random initializations to ensure robustness. Detailed parameters are described in each section before the corresponding experiment results.
IV Results
In this section, we first present the convergence of the proposed model during training in section A. Sections B and C respectively demonstrate the accuracy improvement brought by the proposed joint moment estimation method and the computation speed improvement achieved by using a GRU-based estimator. In section D, we further show the generalization performance of GMF.
IV-A Convergence of Neural Network
@@
Figure 5: (a) Comparison of the accuracy of joint moment estimation for all subjects at different speed ranges using the proposed method, along with the baseline methods. The proposed method improves the prediction accuracy of the model in most cases, especially at faster speeds. (b) Comparison of the overall average and variance of the test results for the three methods.
@@
The convergence of the proposed neural network trained with 10 subjects is illustrated in Fig. 4 (a), represented by the root mean square error (RMSE):
@@
where $\hat{y}_{i}$ denotes the estimated value by the neural network, $y_{i}$ is the ground truth value, and $n$ is the number of samples. During the training process, both the joint moment prediction accuracy and the decoding accuracy of the decoder consistently improved without significant fluctuations. The RMSE of the joint moment estimation on the training dataset, averaged over five repetitions, was 0.0630 ± 0.0014 Nm/kg, while the RMSE of GMF decoding was 0.0016 ± 0.0017 Nm/kg.
IV-B Estimation Accuracy of Models
We evaluated the performance of the proposed method by computing the RMSE of the predicted joint moment on the test dataset. To better demonstrate the effectiveness of the proposed method, we compared our method with two baseline methods: 1) without body parameters, which means directly estimating moment by inputting joint kinematics into the estimator; 2) conventional body parameters fusion model without GMF, which means implementing a fusion model where two networks separately take joint kinematics and body parameters as inputs to obtain two hidden features, and these two features are then stacked and passed through a fully connected network to estimate joint moment. To ensure comparability, the architecture of the estimator inputting joint kinematics is the same for all three baseline models, and the optimizer (Adam) parameters and initial learning rate ($lr=0.001$) are the same. In our method, we set the parameters $w_{1}=1$ and $w_{2}=0.05$ in (17).
Fig. 5 reports the RMSE results of our method compared to the baseline methods. Overall, our method achieves an RMSE of 0.1180 ± 0.0021 Nm/kg on the test dataset. The mean results show a significant improvement compared to the two baseline methods (6.50% vs. baseline method 1 and 8.31% vs. baseline method 2). The experimental results demonstrate that, except for subject 6, whose average joint moment estimation error increased by 0.97% compared to baseline method 1 and 2.57% compared to baseline method 2, the joint moment estimations for other subjects are more accurate. The baseline method 2, which directly fuses body parameters ($\boldsymbol{Q}$) and joint kinematics ($\boldsymbol{X}$) through the network, does not achieve improved joint moment estimation accuracy compared to baseline method 1 with only inputting joint kinematics. In fact, there is even a decrease, indicating that it is challenging to achieve body parameter fusion using conventional fusion models.
As it is difficult to achieve the optimal validation set accuracy with limited time, in addition to using the RMSE, we used the rate of increase in the RMSE on the test set relative to the RMSE on the validation set ($R_{tv}$), which represents the degree of overfitting of the optimal model, obtained from:
@@ IV-B Estimation Accuracy of Models
Fig.4 (b) shows the variation of validation accuracy during the training process for our method and the two baseline methods. All three models achieve similar accuracy on the validation set (around 0.0632 Nm/kg). Based on the statistical analysis (Table. I), our method reduces $R_{tv}$ by 20.06% and 19.20%. While RMSE intuitively demonstrates the effectiveness of our method, the $R_{tv}$ for result evaluation reveals the underlying reason for the accuracy improvement of our method: the mapping of GMF and joint kinematics is more accurate, reducing the prediction inaccuracy of the model in new subject environments.
TABLE I: $R_{tv}$ of proposed method compared to 2 baseline method
Ours Baseline 1 Baseline 2
@@
A complete hip joint moment curve for one gait cycle of a subject at 14 different walking speeds is shown in Fig. 6 (b) for qualitative comparison between the ground truth and predicted values. Furthermore, we analyzed the variation of model estimation error across 28 walking speeds ranging from 0.5 m/s to 1.85 m/s, as depicted in Fig. 6 (a). The estimation error exhibits a decreasing and then increasing trend. For low walking speeds, this could be attributed to the greater variability in gait patterns during slowly walking [33]. On the other hand, the increase in estimation error for fast walking speeds is likely driven by the higher magnitude of joint moments. If statistical analysis is performed using normalized metrics such as normalized root mean squared error (NRMSE), it can be observed that the estimation errors for high walking speeds do not always increase but even decrease compared to moderate walking speeds.
Figure 7: Four joint moment prediction models consisting of estimators with different neural network structures (left), including the proposed GRU-based method for fast estimation and three mainstream model architectures. Average computation time required for a single calculation for each model (right). The blue bar represents the computation time for the GMF estimator, while the red bar represents the total computation time for the joint moment estimation model, including the GMF estimator and decoder.
IV-C Computation Time of Models
To evaluate the performance of the proposed GRU-based estimator in terms of computation time, experiments were conducted on it as well as four existing mainstream network structures(as shown in Fig. 7) for 20 rounds of 10,000 joint moment estimations. All models maintained the structure of the estimator-decoder in the prediction part of Fig. 1.
Previous studies such as [23, 26, 27] have demonstrated that increasing the number of network parameters has a positive impact on the prediction accuracy of joint moment estimation. Therefore, to ensure comparability of the experimental results, we adjusted the network parameter number of the four models to make them as close as possible to the GRU-based model in terms of validation accuracy. Due to device and time limitations, we found it difficult to adjust the TCN to the ideal accuracy, and its optimal validation accuracy was slightly higher than the other networks, indicating that its parameter size needs to be further increased. Therefore, the joint moment estimation time of the TCN network under this experimental condition may be longer. The network parameters used in the experiments are shown in Table. II.
TABLE II: Net Structure Parameters Description
GRU-based Hidden num: 16
Layer num: 1
Full Connected Layer: 16,1
FC-based Layer 1 node num: 32
Layer 2 node num: 16
Layer 3 node num: 16
CNN-based Conv1: kernel 3*1, padding 1, channel 32
Conv2: kernel 3*1, padding 1, channel 64
Conv3: kernel 3*1, padding 1, channel 32
Adaptive Average Pooling: output size 1
Full Connected Layer: 32,1
TCN-based Residual block num: 2
Conv1 in block: kernel 3*1, padding 1, dilation 1
Conv2 in block: kernel 3*1, padding 2, dilation 2
Fig. 7 depicts the computation time of joint moment estimation models based on four estimators. When the input length is 100 time steps, the model with GRU-based estimator demonstrates significantly higher computational efficiency compared to the others. In comparison to estimators with FFN, CNN, and TCN, the computation speed of joint moment estimation is improved by 28.34%, 57.46%, and 58.39% respectively. This result indicates that in scenarios where real-time requirements need to be satisfied, such as in exoskeleton-assisted applications, the model structure based on recurrent neural networks like GRU-based networks have greater potential.
IV-D Generalization of GMF
To verify the generalization of the proposed GMF, the three models with different estimators mentioned in computation time comparison experiment were further used to observe the changes in joint moment estimation accuracy after introducing GMF, following the experimental process outlined in section III-B.
A pre-trained joint moment estimation model with GRU-based estimator was selected for this experimentation. Based on the selected model, we replaced the GRU-based estimator with three estimators with different network structures. Following the training process shown in Fig. 1, a series of new training processes were conducted. During the training process, only the parameters of the new estimator were updated while the parameters of the generator and decoder were fixed (keeping the generated GMF unchanged).
TABLE III: Comparison of the prediction accuracy of the model with the GMF and the models without the GMF
Model Model without body parameters Fusion model without GMF Ours

### v2

@@ 2.2.3 GMF Estimator
In exoskeleton systems, recurrent neural networks (RNN) can better leverage their rapid computation ability [28] to estimate joint moments. Existing recurrent neural network structures include RNN, LSTM, GRU [29]. Due to the challenges of RNN in handling long-term dependencies, vanishing gradients, and exploding gradients, they have been gradually replaced by LSTM and GRU. Furthermore, GRU has a more simplified structure and fewer parameters compared to LSTM, resulting in lower memory usage and faster computation efficiency. In many cases, there is no significant difference in estimation accuracy between the two models [30]. Therefore, the estimator in this study utilizes a neural network based on the GRU structure.
@@
In this study, only one GRU block is used as an implementation case, with 16 hidden neurons, and a fully connected network with a size of 16 $\times$ 1 is used to output the predicted target (Fig. 2). Specifically, at each time step, the GRU receives two different modalities: the hidden information from the previous time step’s GRU hidden layer and the inputs $\boldsymbol{X}$. It adjusts the information flow through update gates and reset gates, and finally obtains the output based on these two gates. Thus, given the hip joint sagittal plane angle signal at time t as $\boldsymbol{X}_{t}$, the hidden inputs from the previous time step as $\boldsymbol{H}_{t-1}$, the estimated GMF at the current time step $\widehat{GMF}$ can be obtained as follows:
$\boldsymbol{H}_{t}=\text{GUR}(\boldsymbol{X}_{t},\boldsymbol{H}_{t-1}),$ (6)
$\widehat{GMF}=\text{FC}(\boldsymbol{H}_{t}).$ (7)
Writing the network parameters ($\boldsymbol{W}$ and $\boldsymbol{b}$) as $\boldsymbol{\theta}_{E}$, like (5), the optimal parameters $\boldsymbol{\theta}_{E}^{*}$ are solved as:
$\boldsymbol{\theta}_{E}^{*}=\mathop{\textbf{argmin}}\limits_{\boldsymbol{\theta}_{E}}\left\|E(\boldsymbol{X};\boldsymbol{\theta}_{E})-G(\boldsymbol{q},M;\boldsymbol{\theta}_{G})\right\|_{2}.$ (8)
@@
2.2.4 GMF Decoder
The GMF decoder is used to decode the generated GMF into joint moment. We employ the same network structure as the GMF generator for constructing the GMF decoder. To improve the accuracy of the output, the body parameters are included as part of its inputs. Using $\boldsymbol{\theta}_{D}$ to represent the network parameters of the GMF decoder, the optimal parameters $\boldsymbol{\theta}_{D}^{*}$ of the GMF decoder in the above process can be obtained from:
$\boldsymbol{\theta}_{D}^{*}=\mathop{\textbf{argmin}}\limits_{\boldsymbol{\theta}_{D}}\left\|D(\boldsymbol{q},GMF;\boldsymbol{\theta}_{D})-M\right\|_{2}.$ (9)
2.3 Model Training
The proposed method consists of two parts: GMF estimation and GMF decoding, which involves the training of three models: estimator, generator, and decoder. In the first part, the GMF generator is jointly trained with the estimator to obtain better individual-independent representation (GMF) and its estimation model (estimator). Meanwhile, in the second part, the GMF generator is trained together with the decoder to ensure the decodability of the generalized GMF and to obtain a model capable of decoding the GMF into the joint moment.
Based on (5), (8), and (9), two loss functions $L_{1}$, $L_{2}$ are constructed for the network training. The backpropagation process is shown in Fig. 1 - Model training with solid lines. $L_{1}$ represents the loss between the results of the GMF estimator and the GMF generator, while $L_{2}$ represents the loss between the outputs of the GMF decoder and ground-truth values of the joint moment. The loss functions are defined as follows:
$J_{mse}(y_{i},\hat{y_{i}})=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y_{i}})^{2},$ (10)
$L_{1}=J_{mse}(F(\boldsymbol{X}),G(\boldsymbol{q},M)),$ (11)
$L_{2}=J_{mse}(D(G(\boldsymbol{q},M),\boldsymbol{q}),M).$ (12)
It is worth noting that the optimization of the parameters of the GMF generator is also included in $L_{2}$. By selecting the weighting coefficients $w_{1}$ and $w_{2}$ of the two loss functions, it can be ensured that the GMF, which is generated by the feature generator and optimized by $L_{1}$, can be decoded into the joint moment through the GMF decoder. Therefore, the condition in (5) is satisfied. The overall loss function of the joint moment estimation network is given by:
@@ 2.3 Model Training
2.4 Model Deployment
The trained GMF estimator and decoder will be used in the estimation process for the final joint moment estimation. As shown in Fig. 1 - Model deployment, using the joint kinematics as input, the GMF is obtained by the fast GMF estimator, which is then combined with user body parameters to be decoded into the predicted joint moment:
$\widehat{M}=D(E(\boldsymbol{X}),\boldsymbol{q}).$ (14)
Subsequently, the estimated joint moments are converted into assistive torques by the mid-level and low-level controllers. In this study, we implement a three-step middle-level exoskeleton controller [11, 31] to provide users assistance during walking.
Figure 3: Exoskeleton and controller used for validating the proposed method. Motors with encoder provide the controller hip joint angles as inputs and generate assistance torque to assist users. The controller is a three-level structure, where the high-level estimates hip joint moments of the user, the mid-level processes hip joint moments to obtain the desired assistance torque and the low-level FOC generates motor torque stably.
A hip exoskeleton prototype was designed to implement assistance experience for users. As shown in Fig. 3, the assistance torque is obtained by scaling, delaying, and filtering the hip moment estimates. Specifically, $\tau_{\text{assist}}$ at time $t$ is calculated by:
$\tau_{\text{assist},t}=\text{Low-pass Filter}(k\cdot\widehat{M}_{t-\delta}),$ (15)
where $k$ is the scale factor and $\delta$ is the delay factor.
3 Experiments
There were two designed experiment parts in this study: the first part was to validate the GMF-based model by a dataset, which is aimed at evaluating the accuracy and generalization of the proposed method, and the second part was to apply the trained model on a hip exoskeleton to validate its performance (assistance gain for users) when used as an exoskeleton high-level controller.
This work involved human subjects in its research. Approval of all ethical and experiment procedures and protocols was granted by Medical Ethics Committee of Southern University of Science and Technology [No.20240022], under Application 2024-02-24. All participants provided written informed consent to participate in the study.
3.1 Implementation Details
We conducted model training on a device with a CPU (Intel, Core i7-12700F) and an GPU (Nvidia, GeForce RTX 3060Ti). The Adam optimizer [32] was employed. In all experiments, we performed five repetitions with different random initializations to ensure robustness. Detailed parameters are described in each section before the corresponding experiment results. The model deployment (exoskeleton controller) was implemented on a Mini-PC (Intel, NUC11PH).
3.2 Validation on dataset
3.2.1 Dataset and Training-testing Data Partitioning
@@
3.2.2 Models for accuracy validation
We evaluated the performance of the proposed method by computing the RMSE of the predicted joint moment on the test dataset. To better demonstrate the effectiveness of the proposed method, we compared our method with two baseline methods:
  • 1.
    Model without body parameters - means directly estimating moment by inputting joint kinematics into the estimator without body parameters.
  • 2.
    Fusion model without GMF - means implementing a conventional fusion model where two networks separately take joint kinematics and body parameters as inputs to obtain two hidden features, and these two features are then stacked and passed through a fully connected network to estimate joint moment.
To ensure comparability, the architecture of the part inputting joint kinematics in two baseline neural networks is the same with the GMF estimator, and the optimizer (Adam) parameters and initial learning rate ($lr=0.001$) are the same. In our method, we set the parameters $w_{1}=1$ and $w_{2}=0.05$ in (13). The baseline models are represented as:
$\widehat{M}_{\text{b1}}=E_{\text{b1}}(\boldsymbol{X}),$ (16)
$\widehat{M}_{\text{b2}}=FC_{\text{b2}}(E_{\text{b2}}(\boldsymbol{X}),\boldsymbol{q}).$ (17)
3.2.3 Generalization validation
Figure 5: Four joint moment estimation models consisting of estimators with different neural network structures, including the proposed GRU-based method for fast estimation and three mainstream model architectures.
More neural network structures were used to verify the proposed GMF framework. To ensure the generalization of the proposed GMF, three existing mainstream network structures (FCNN, CNN, TCN) were further used to observe the changes in joint moment estimation accuracy after introducing GMF as shown in Fig. 5, following the experimental process outlined in 3.2.2.
  • 1.
    The FCNN - was composed of three fully connected layers. The first layer contained 32 nodes, followed by two hidden layers with 16 nodes each.
  • 2.
    The CNN - consisted of three convolutional layers, each with a 3×1 kernel size and padding of 1. The layers progressively processed features with 32, 64, and 32 channels, respectively. The extracted features were then passed through an adaptive average pooling layer with an output size of 1, followed by a fully connected layer with 32 and 1 neurons.
  • 3.
    The TCN - employed two residual blocks, each containing two convolutional layers. The first convolutional layer within each block had a 3×1 kernel size, 32 channels, padding of 1, and a dilation of 1. The second convolutional layer maintained the same kernel size and number of channels but applied a dilation of 2 with padding of 2.
To ensure that the variables are consistent, all these models utilized the same GMF generator and decoder, which are trained with the GRU-based estimator in 3.2.2. Specifically, we replaced the GRU-based estimator with the three different network structures. Following the training process shown in Fig. 1, a series of new training processes were conducted. During the training process, only the parameters of the new estimator were updated while the parameters of the generator and decoder were fixed (keeping the generated GMF unchanged).
Previous studies such as [23, 26, 27] have demonstrated that increasing the number of network parameters has a positive impact on the estimation accuracy of joint moment estimation. Therefore, to ensure comparability of the experimental results, we adjusted the network parameter number of the three models to make them as close as possible to the GRU-based model in terms of validation accuracy when conducting the first baseline as (16).
3.3 Assistance for Human Walking
The poor generalized model might lead to the controller instability, which reduces the exoskeleton performance and even results in a fall for the user [11]. Therefore, real-world validation is critical for instantaneous joint moment estimation methods used for exoskeleton control. We validated the proposed method on an exoskeleton and implemented walking experiments with users wearing it.
Considering the validation effect, we adopted a rigid exoskeleton structure, ensuring the full torque transfer from the motors to the user as possible. The candidate coefficients of (15) were set as $k=0.15$ and $\delta=0.3$ s, which were determined through preliminary experiments.
To quantify the gain for reducing efforts when users walk with the proposed high-level controller, comparison experiments were conducted. Five participants (age 23.4 ± 1.4 y, height 1.7 ± 0.2 m, mass 60.6 ± 1.4 kg) were recruited to complete metabolic trials under three conditions: No Exo, Assist off, and Assist On. The K5 portable metabolic system (COSMED, Rome, Italy) was used to record the experiment data. Specifically, A subject took part in 3 trials with their metabolic cost being collected: a trial without the hip exoskeleton including neutral standing for 3 minutes and walking at 1.25 m/s for 5 minutes, a trial with the no-powered exoskeleton for 5 minutes walking, and a trial with the powered exoskeleton for 5 minutes walking. At least a 10-minute break between trials to allow subjects to return to normal metabolic levels. Before each trial, the participant was required to fast for a minimum of 2 h to reduce metabolic bias from food intake. The Instantaneous metabolic cost $E$ was computed based on $\dot{V}O_{2}$ and $\dot{V}CO_{2}$ (collected by K5) as:
$E=\frac{16.48*\dot{V}O_{2}+4.48*\dot{V}CO_{2}}{60*m},$ (18)
where $m$ is the body mass of the subject, and the net metabolism rate $E_{\text{net}}$ was obtained by:
$E_{\text{net}}=E_{\text{walking}}-E_{\text{static}}.$ (19)
4 Results
In this section, the results of the experiments are presented with detailed statistical analysis. The convergence of the proposed model during training is illustrated in Subsection 4.1. Subsections 4.2 and 4.3 demonstrate the accuracy improvement brought by the proposed GMF method and the generalization performance of GMF. Subsection 4.4 shows the computation efficiency advantage of using a GRU-based estimator on the exoskeleton controller. In Subsection 4.5, the tested assistance gain for users with the proposed method is given.
Figure 6: (a) Trend of the accuracy of the joint moment estimation (dark blue) and accuracy of GMF decoding (light blue) of training data during the training process of the proposed estimation model based on GMF. Both values of them continuously improve and converge to the optimum over the 5000 epochs. (b) The trend of estimation accuracy on the validation set for our method (dark red) compared to two other contrastive methods (light red). The three methods approximate the best accuracy achieved under the experimental conditions set.
@@ 4 Results
Figure 7: (a) Comparison of the accuracy of joint moment estimation for all subjects at different speed ranges using the proposed method, along with the baseline methods. The proposed method improves the estimation accuracy of the model in most cases, especially at faster speeds. (b) Comparison of the overall average and variance of the test results for the three methods.
4.1 Convergence of Neural Network
@@
The convergence of the proposed neural network trained with 10 subjects from dataset is illustrated in Fig. 6 (a), represented by the root mean square error (RMSE):
@@
where $\hat{y}_{i}$ denotes the estimated value by the neural network, $y_{i}$ is the ground truth value, and $n$ is the number of samples. During the training process, both the joint moment estimation accuracy and the decoding accuracy of the decoder consistently improved without significant fluctuations. The RMSE of the joint moment estimation on the training dataset, averaged over five repetitions, was 0.0630 ± 0.0014 Nm/kg, while the RMSE of GMF decoding was 0.0016 ± 0.0017 Nm/kg.
4.2 Estimation Accuracy of Models
Fig. 7 reports the RMSE results of our method compared to the baseline methods. Overall, our method achieves an RMSE of 0.1180 ± 0.0021 Nm/kg on the test dataset. The mean results show a significant improvement compared to the two baseline methods (6.50% vs. model without body parameters and 8.31% vs. fusion model without GMF). The experimental results demonstrate that, except for subject 6, whose average joint moment estimation error of all walking speed increased by 0.97% compared to model without body parameters and 2.57% compared to fusion model without GMF, the joint moment estimations for other subjects are more accurate. The second baseline method, which directly fuses body parameters ($\boldsymbol{q}$) and joint kinematics ($\boldsymbol{X}$) through the network, does not achieve improved joint moment estimation accuracy compared to model without body parameters with only inputting joint kinematics. In fact, its performance slightly degrades, suggesting that conventional fusion models struggle to effectively integrate body parameters.
As it is difficult to achieve the optimal validation set accuracy with limited time, in addition to using the RMSE, we used the rate of increase in the RMSE on the test set relative to the RMSE on the validation set, denoted as $R_{tv}$, which represents the degree of overfitting [34, 35] of the optimal model, obtained from:
@@ 4.2 Estimation Accuracy of Models
Fig.6 (b) shows the variation of validation accuracy during the training process for our method and the two baseline methods. All three models achieve similar accuracy on the validation set (around 0.0632 Nm/kg). Based on the statistical analysis (Table. 1), our method reduces $R_{tv}$ by 20.06% and 19.20%. While RMSE intuitively demonstrates the effectiveness of our method, the $R_{tv}$ for result evaluation reveals the underlying reason for the accuracy improvement of our method: the mapping of GMF and joint kinematics is more accurate, reducing the estimation inaccuracy of the model in new subject environments.
Table 1: $R_{tv}$ of proposed method compared to baseline methods
Model with GMF Model without body parameters Fusion model without GMF
@@
Table 2: Estimation accuracy performance with different estimators in the generalization experiment
Estimator Model without body parameters Fusion model without GMF GMF-based model (Ours)
@@
Complete hip joint moment curves for one gait cycle of a subject at 14 different walking speeds is shown in Fig. 8 (b) for qualitative comparison between the ground truth and predicted values. Furthermore, we analyzed the variation of model estimation error across 28 walking speeds ranging from 0.5 m/s to 1.85 m/s, as depicted in Fig. 8 (a). The estimation error exhibits a decreasing and then increasing trend. For low walking speeds, this could be attributed to the greater variability in gait patterns during slow walking [36]. On the other hand, the increase in estimation error for fast walking speeds is likely driven by the higher magnitude of joint moments, as shown in Fig. 8 (b).
4.3 Generalization of GMF